# MetaAligner: Conditional Weak-to-Strong Correction for Generalizable Multi-Objective Alignment of Language Models

### v1

@@ 3.4 Generalizable Inference
Table 2: Performance of MetaAligner-(1.1B, 7B, 13B) on each IMHI sub-task over different policy models. The GPT-4 judge considers 3 objectives: Correctness, Informativeness, and Professionality. The figures show the advantage of aligned outputs over the policy model outputs on win rate. Best values for each MetaAligner model are highlighted in bold.
4 Experiments
4.1 Experimental Settings
Datasets
We train and evaluate the performance of MetaAligner on the IMHI dataset (Yang et al., 2023b). For the training data, we invite domain experts to manually label 7.2K response pairs sampled from the training split of IMHI, considering 3 objectives with the following text markers: (1) Correctness: the response should make correct predictions in the corresponding sub-tasks; (2) informativeness: the response should express clear logic and provide consistent supporting evidence; (3) Professionality: the generated responses should provide supportive evidence with high quality and reliability. For evaluation, we perform alignment on the IMHI benchmark, which consists of 9 sub-tasks with 20K query-response pairs.
Models
We train MetaAligner-(1.1B, 7B, 13B) models based on TinyLLaMA-1.1B (Zhang et al., 2024) and LLaMA2-Chat-(7B, 13B) (Touvron et al., 2023) foundation models. We utilize MetaAligner to perform multi-objective alignment on the following policy models: LLaMA2-Chat-(7B,13B,70B) (Touvron et al., 2023), Gemma-instruct-(2B,7B) (Team et al., 2024), MentaLLaMA-(7B, 13B, 33B) (Yang et al., 2023b), and Vicuna-(7B, 13B, 33B) (Chiang et al., 2023). Among the policy models, MentaLLaMA models are instruction-tuned on mental health-related data, and other models are tuned on general-domain instruction data.
Evaluation Metric
We quantify the alignment performance of model outputs by comparing their win rate against the ground-truth response provided by the IMHI benchmark, on the corresponding objectives. Considering the large amounts of test samples, we utilize GPT-4, a widely utilized evaluation tool in previous works (Guo et al., 2024; Sun et al., 2023; Li et al., 2023a), to perform the judgments, where each target response, query, ground-truth response, and objectives are provided via prompt engineering. GPT-4 is required to select the response with higher alignment with the specified objectives in the prompt, or indicate a tied performance.
Figure 2: Objective-wise win-rate comparisons based on different policy models. MetaAligner-7B with different levels of alignment are tested on Correctness, Informativeness, and Professionality, compared with 3 different policy models. The win rates are obtained via the judgments of GPT-4 on the corresponding objectives.
4.2 Overall Performance
We stack MetaAligner on different policy models to perform alignment on all 3 objectives: Correctness, Informativeness, and Professionality. The overall performance on the IMHI benchmark and its separation into 9 different sub-tasks are shown in Table 2.
@@ 4.2 Overall Performance
Within most policy model families, we observe a decrease in win-rate advantage as their sizes increase. These decreases indicate a struggle aligning powerful large-scale models with a small MetaAligner. Fortunately, MetaAligner’s capabilities also show scalability. Increasing the size of its base model leads to a higher win-rate advantage on most policy models. Specifically, MetaAligner-7B outperforms or achieves comparable performance to MetaAligner-1.1B on 9 policy models, and MetaAligner-13B surpasses MetaAligner-7B on 10 policy models by an average of 12.58%. These observations motivate further explorations in model scale-performance balance for MetaAligner.
Figure 3: Generalizable alignment on 3 unseen objectives based on LLaMA2-Chat-70B outputs. In the x-axis, "+" denotes the addition of a new objective into the objective set $\mathcal{O}$. "★" denotes the unaligned win rates for LLaMA2-Chat-70B outputs. "-." lines identify win rate fluctuations before alignment on the corresponding objective, and solid lines identify fluctuations after alignment.
4.3 Objective-Wise Evaluation
We evaluate the objective-wise performance of MetaAligner by decoupling the target objectives. We utilize MetaAligner to perform 5 levels of alignments: No alignment, aligning on each objective (Correctness, Informativeness, and Professionality), and full alignment. The objective-wise win rates are shown in Figure 2. We have the following observations:
Objective-wise alignment significantly improves the target objective and benefits the performance on other objectives. For example, aligning on Informativeness with MetaAligner achieves the best performance on Informativeness, and increases win rates on Correctness and Professionality over the unaligned outputs. This tendency holds with all policy models and other objectives. These results further prove the complex interplay within multi-objective alignment, where positive correlations also exist, in addition to contradictions (Guo et al., 2024).
Aligning on all objectives provides balanced performance. According to the results, full alignment displays competitive performance on all 3 objectives and generally outperforms unaligned outputs and aligned outputs from other objectives, even comparable to aligned outputs from the same objective in some cases. The reason is that MetaAligner learns weak-to-strong corrections based on dynamic objective conditions, enabling the model to fully attend to the specified objectives and achieve a Pareto optimal correction on these conditions.
MetaAligner performs comprehensive alignment on different sub-tasks. With most policy models, MetaAligner significantly outperforms the unaligned responses on all 9 sub-tasks, regardless of the alignment objectives. These results indicate the general effectiveness of our dynamic-objective method on tasks demanding various expertise (Yang et al., 2023b). However, MetaAligner still struggles to align certain sub-tasks, especially with strong policy models. For example, MetaAligner outputs fail to outperform MentaLLaMA-13B on the CAMS sub-task in all objectives. These results show enhancements in domain knowledge, with techniques such as retrieval-augment generation (Lewis et al., 2020), a promising future direction for improvement.
Algorithm Space Complexity Correctness Informativeness Professionality Overall
MORLHF 22.27$\times$ -6.31% +33.7% +16.31% +15.08%
MODPO 17.27$\times$ +26.31% +30.1% +33.12% +32.9%
SFT 11.82$\times$ +19.22% +24.89% +25.6% +26.62
MetaAligner-1.1B 1$\times$ +26.11% +35.67% +13.11% +29.7%
MetaAligner-7B 6.36$\times$ +28.89% +24.92% +29.07% +27.37%
MetaAligner-13B 11.82$\times$ +31.33% +45.0% +43.66% +35.55%
Table 3: Win rates of previous alignment methods and MetaAligner models based on the LLaMA2-Chat-13B policy model. Space complexity is calculated based on the memory consumption of training MetaAligner-1.1B.
4.4 Generalizable Alignment
In this section, we explore zero-shot preference alignment by utilizing MetaAligner to align with 3 less relevant objectives that are unseen during its training phases: (1) Completeness (Yang et al., 2023a): the response should cover all relevant aspects of the original query; (2) Honesty (Guo et al., 2024): the response should speak in truth that is consistent with the query and commonsense. (3) Safety (Sun et al., 2023): The response should avoid content that is offensive, discriminatory, or harmful.
2,700 random-selected responses from the IMHI benchmark are re-aligned with these unseen objectives added to the objective set $\mathcal{O}$ one-by-one. Their performance is presented in Figure 3. According to the results, we have the following conclusions:
MetaAligner performs effective zero-shot alignment for unseen objectives. With all MetaAligner models, incorporating an unseen objective significantly improves its corresponding win rate. For example, MetaAligner-7B improves win rate by 40.7% on Completeness, 28.89% on Honesty, and 17.56% on Safety compared to each of these objectives not introduced. These results prove the viability of generalizable alignment with the in-context learning ability of MetaAligner. In addition, alignment performance on supervised objectives (Correctness, Informativeness, Professionality) generally surpasses unseen objectives on all tested models, showing that MetaAligner also learns objective-specific features during training.
Performance on aligned objectives is maintained with more unseen alignment objectives. As each objective is aligned, its win rate surges, then stabilizes as long as the objective is included in $\mathcal{O}$. These results prove MetaAligner to perform reliable alignment with the expansion of objectives. However, aligning on new objectives can affect other unaligned objectives. For example, aligning on Correctness with MetaAligner-1.1B benefits its win rate by 11.22%, but harms performance on 3 other unaligned objectives. These problems are due to the limited ability of smaller MetaAligner models and the controversial dependence of certain multi-objective alignments (Guo et al., 2024), where we observe an alleviation as the model size increases.
MetaAligner’s generalizability also shows scalability. Performance on the 3 unseen objectives increases with the scale-up of MetaAligner model size, where MetaAligner-1.1B improves 12.7% on average, MetaAligner-7B improves 29.05%, and MetaAligner-13B improves 36.69%. This scalability is attributed to the growing in-context learning ability of larger base models, which motivates further explorations into the correlation between generalizable alignment and model scales.
4.5 MetaAligner vs. MORLHF/MODPO/SFT
We compare the performance of MetaAligner with MORLHF, MODPO, and SFT-based methods on 2700 random-selected responses from the IMHI benchmark. We implement the linear scalarization method for MORLHF. We follow previous settings of realizing MODPO (Guo et al., 2024) by pre-pending the objective markers. Due to the closed-source nature of current SFT-based methods, we only replicate stage 1 where details are clear, and utilize the preference-to-reward mapping method proposed in Yang et al. (2024). We select LLaMA2-Chat-13B as the policy model, and the results are presented in Table 3.
According to the results, MetaAligner-13B achieves over 30% of win rates on all objectives, and generally outperforms all previous methods by 10.68% on average. This advantage is achieved with equal or less space complexity than previous algorithms, showing the high efficiency of the proposed method. As the base model size further reduces, MetaAligner-1.1B and MetaAligner-7B show lower, but still competitive performance to the 13B model. Impressively, MetaAligner-1.1B outperforms both MORLHF and SFT-based methods by 8.85% on average, and approaches MODPO with merely a 3.2% gap, but with 11.82$\times$-22.27$\times$ less computational resources. This huge advantage indicates a high prospect for applying MetaAligner in low-resource scenarios. For previous methods, MODPO and SFT can provide stable alignment for multiple objectives, but with high space complexity. MORLHF bears the highest resource consumption, but achieves the lowest 15.08% overall win rate. Its training process is also unstable, including a 6.31% decrease in the Correctness objective.
5 Related Work
5.1 Large Language Models
Large language models (LLMs) have reached near-human capabilities across a wide spectrum of tasks related to understanding, generating, and reasoning with natural language Achiam et al. (2023); Team et al. (2023); Touvron et al. (2023); Luo et al. (2023). Notable examples include commercially available LLMs like ChatGPT, GPT-4 Achiam et al. (2023), and Gemini Team et al. (2023), as well as leading-edge open-source models like LLaMA2 Touvron et al. (2023) and Mistral Jiang et al. (2023). Despite their advancements, these LLMs can exhibit problematic behaviors, including the generation of inaccurate information Zhang et al. (2023); Xie et al. (2023), flattery, and deception, raising concerns about their potential negative impacts on society and associated risks Bommasani et al. (2021). To address these issues, considerable research efforts have been dedicated to refining LLMs’ outputs to better align with human values and preferences, employing methods such as supervised fine-tuning and reinforcement learning Ji et al. (2023).
5.2 Preference Alignment
Many studies, featuring leading large language models such as ChatGPT, GPT-4, and LLaMA2, have delved into enhancing characteristics like helpfulness, harmlessness, and honesty. These efforts have largely centered on Reinforcement Learning from Human Feedback (RLHF) Ouyang et al. (2022). Nonetheless, employing a singular reward model proves inadequate for aligning LLMs with varied human preferences. This limitation has motivated the exploration of multi-objective RLHF (MORLHF) Sener and Koltun (2018); Li et al. (2020); Rame et al. (2024), by linear scalarizations of multiple rewards Sener and Koltun (2018); Li et al. (2020) and interpolations of LLM weights trained from diverse reward models Rame et al. (2024). However, diverse reward models can increase the computational cost and lead to challenges of training due to the unstable nature of the PPO, some studies further explore the multi-objective direct preference optimization (MODPO) Zhou et al. (2023) without the RL paradigm. Zhou et al. (2023) proposed multi-objective direct preference optimization (MODPO), which extended DPO algorithm (Rafailov et al., 2024) to combine multiple objectives with specific weightings. Another paradigm is the SFT-based methods Yang et al. (2024); Guo et al. (2024), which use customized prompting strategies to explicitly incorporate multiple reward values and optimize with supervised fine-tuning (SFT) or DPO. These methods also facilitated objective-wise controllable generation during inference.
6 Discussion
Conclusion.
This paper proposed MetaAligner, the first policy-agnostic and generalizable method for multi-objective preference alignment. Based on the dynamic multi-objective dataset, It follows a three-step training paradigm to obtain flexible alignment ability, which enables plug-and-play inference and zero-shot expansion to unseen objectives. Thorough investigations on LLMs ranging from 2B to 70B proved MetaAligner’s overall effectiveness in multi-objective alignment for various policy models and objective-wise alignment. and Further experiments showed its strong generalizability to unseen objectives and scalability to simultaneously align more objectives.
Limitation.
Though achieving outstanding performance, MetaAligner still bears limitations. Firstly, similar to the vanilla Aligner module Ji et al. (2024), stacking MetaAligner module on policy models inevitably leads to increased computational burdens during alignment inference, which affects actual model deployment, especially for scenarios such as local deployment on mobile devices. Secondly, MetaAligner still provides limited improvement for large-scale domain-specific policy models, such as MentaLLaMA-33B, in domain knowledge-intense tasks such as mental health analysis. These results reflect the importance of expertise knowledge for multi-objective alignment, where MetaAligner still requires further enhancements. Thirdly, due to limited resources, we only tested the generalizability of MetaAligner on 3 unseen objectives, which does not provide a clear landscape of its alignment performance on more objectives.
Future Work.
Firstly, we aim to explore improving MetaAligner in domain-specific alignment scenarios utilizing techniques such as retrieval-augment generation (Lewis et al., 2020). Secondly, we will dive deep into the scalability of MetaAligner to evaluate its impact on alignment performance, including the model scale-performance balance. We will also provide a clearer landscape of their generalizable alignment ability by examining larger base model sizes and aligning on much more unseen objectives (we only expanded to 6 objectives). This analysis will be valuable guidance in leveraging MetaAligner for generalizable multi-objective alignment.

### v2

@@ Appendix D Sub-task Performance on IMHI Benchmark
Table 7: Performance of MetaAligner-(1.1B, 7B, 13B) on each IMHI sub-task over different policy models. The GPT-4 judge considers 3 objectives: "Correct", "Informative", and "Professional". The figures show the advantage of aligned outputs over the policy model outputs on win rate. Best values for each MetaAligner model are highlighted in bold.
We stack MetaAligner on different policy models to perform alignment on all 3 objectives: "Correct", "Informative", and "Professional". We include MentaLLaMA-(7B, 13B, 33B) [34], the first open-source instruction-following LLM series for interpretable mental health analysis into the policy models. Details about the 9 sub-tasks are provided in Table 6. The overall performance of MetaAligner on the IMHI benchmark and its separation into 9 different sub-tasks are shown in Table 7.
@@
Appendix E Objective-wise Alignment
E.1 Experimental Settings
We randomly sample 1,200 queries from the UltraFeedback test set and probe the target policy models to provide responses to all queries, which are regarded as unaligned responses. Then MetaAligner is used to align these responses under different objectives, including aligning on each objective (Instruction following, Helpfulness, Honesty, and Truthfulness), and another full alignment on all 4 objectives. During evaluation, GPT-4 is leveraged as a reward model to score these responses considering different objectives. Specifically, the following prompt is developed to obtain the scores:
[$\mathcal{S}(q,a,\left<r\right>)$] You are a skilled evaluator of helpful AI assistants. You will be presented with one query and a response to this query. QUERY: {q} | RESPONSE: {a} Assign a score ranging from 0 to 10 to this response considering the following aspect: $r\right>$}. The rating improves as the score rises, where 0 denotes inferior performance and 10 denotes approaching-human performance. You must only return the assigned score.
where $q$ and $a$ denotes the target query and response, and $r\in${Instruction following, Helpful, Honest, Truthful} is the target objective for evaluation.
E.2 Experimental Results
More experimental results are presented in Figure 4. The results are the performance of MetaAligner-7B on Gemma-instruct-7B, LLaMA2-Chat-70B, and GPT-3.5 outputs from the UltraFeedback test set. According to the results, MetaAligner can perform effective objective-wise alignment on outputs of different policy models, from the small Gemma-7B model to the competitive commercial models GPT-3.5. Unlike full alignment which shows weaker performance as the capability of the policy model increases, objective-wise alignment provides stable improvement for different policy models. For example, aligning "Instruction following" leads to over 20% of approaching-human responses for all policy models, and aligning "Honesty" leads to over 30% of approaching-human responses for all policy models. The reason is that single-objective alignment doesn’t involve the complex interactions between multiple objectives, which facilitates the full realization of MetaAligner’s potential to improve the response on the corresponding objective.
Figure 4: Objective-wise kernel density estimates of GPT-4 evaluation scores under different alignment objectives.
Appendix F Implementation Details of Baseline Models
F.1 SFT-based Methods
Existing SFT-based methods [11, 35] for multi-objective preference alignment explicitly incorporate reward values into the query via prompt engineering, which includes a text marker for each alignment objective and their corresponding value numbers in the current response. The model is trained to predict the response given the enhanced query, which enables it to learn a mapping between the reward value and its reflection in the response. During inference, we achieve alignment by assigning a higher reward value to the target objectives. Specifically, we define the following prompting template for HH-RLHF dataset:
[$\mathcal{F}(q,r_{1},r_{2},r_{3}$] <Harmlessness>: {$r_{1}$}; <Helpfulness>: {$r_{2}$}; <Humour>: {$r_{3}$} | {q}
where $q$ denotes the query, $r_{1}$, $r_{2}$, $r_{3}$ denote the corresponding reward values for the current response, ranging from 1 to 5. We define the following prompting template for UltraFeedback dataset:
[$\mathcal{F}(q,r_{1},r_{2},r_{3},r_{4}$] <instruction_following>: {$r_{1}$}; <honesty>: {$r_{2}$}; <honesty>: {$r_{3}$}; <helpfulness>: {$r_{4}$} | {q}
where $q$ denotes the query, $r_{1}$, $r_{2}$, $r_{3}$, $r_{4}$ denote the corresponding reward values for the current response, obtained from existing reward models.
During inference, we aim to simultaneously align all objectives with 1 model to enable fair comparisons to MetaAligner. For UltraFeedback, since all rewards range from 1 to 5, we set all values to 5 during inference: $r_{1}=r_{2}=r_{3}=r_{4}=5$. For HH-RLHF, there are no unified ranges for each objective as all training values are obtained from reward models. Therefore, we set each objective value to its maximum in the training dataset to enable higher alignment. The values are set to: $r_{1}=4.19$, $r_{2}=3.03$, $r_{1}=1.58$.
In calculating the GPU hours, we sum the training hours for fine-tuning the LLaMA-Chat-7B policy model for HH-RLHF and UltraFeedback datasets.
F.2 MODPO
We implement MODPO by implementing the Controllable Direct Preference Optimization (CDPO) [11] algorithm based on its paper introduction. We bypass the controllable preference SFT stage by utilizing the trained model from SFT-based methods as the foundation model. In the CDPO stage, each query $q$ is accompanied by two responses $c_{1}$ and $c_{2}$, where each response is constructed into contrastive pairs. In HH-RLHF, we have the following prompting template:
[$\mathcal{F}(q,c_{1},r_{1}^{1},r_{1}^{2},r_{1}^{3}$] <Harmlessness>: {$r_{1}^{1}$}; <Helpfulness>: {$r_{1}^{2}$}; <Humour>: {$r_{1}^{3}$} | {q} | {$c_{1}$}
[$\mathcal{F}(q,c_{2},r_{2}^{1},r_{2}^{2},r_{2}^{3}$] <Harmlessness>: {$r_{2}^{1}$}; <Helpfulness>: {$r_{2}^{2}$}; <Humour>: {$r_{2}^{3}$} | {q} | {$c_{2}$}
where $r_{i}^{j}$ denotes the reward value for the $i$-th response on the $j$-th objective. For HH-RLHF, there are no unified ranges for each objective as all training values are obtained from reward models. Therefore, we set each preference value to its maximum in the training dataset to enable higher alignment. The values are set to: $r^{1}=4.19$, $r^{2}=3.03$, $r^{3}=1.58$. Similarly, on UltraFeedback we have:
[$\mathcal{F}(q,c_{1},r_{1}^{1},r_{1}^{2},r_{1}^{3},r_{1}^{4}$] <instruction_following>: {$r_{1}^{1}$}; <honesty>: {$r_{1}^{2}$}; <honesty>: {$r_{1}^{3}$}; <helpfulness>: {$r_{1}^{4}$} | {q} | {$c_{1}$}
[$\mathcal{F}(q,c_{2},r_{2}^{1},r_{2}^{2},r_{2}^{3},r_{2}^{4}$] <instruction_following>: {$r_{2}^{1}$}; <honesty>: {$r_{2}^{2}$}; <honesty>: {$r_{2}^{3}$}; <helpfulness>: {$r_{2}^{4}$} | {q} | {$c_{2}$}
For UltraFeedback, since all rewards range from 1 to 5, we set all values to 5 during the DPO training process: $r^{1}=r^{2}=r^{3}=r^{4}=5$. In determining the preference scores, we use the CDPO learning goal to transform the task into a conditional multi-objective optimization problem. Specifically, the reward value $R_{i}$ for response $c_{i}$ is calculated as follows:
$R_{i}=\sum_{i=1}^{m}\omega_{i}g_{i}$ (9)
where $\omega_{i}$ represents the weight of the $i$-th objective and $g_{i}$ is calculated as follows:
$\small g_{i}=\begin{cases}-\lambda_{i}\lvert p_{j}-r_{i}^{j}\rvert,&\text{if $i$-th objective is controlled,}\\ r_{i}^{j},&\text{otherwise.}\end{cases}$ (10)
where $\lambda_{i}$ represents the weight of the controlled objective and $p_{j}$ is a pre-defined preference value for the $j$-th objective. In our implementation, we set $\omega_{i}=\lambda_{i}=1$. We aim to simultaneously align all objectives with 1 model to enable fair comparisons to MetaAligner. For UltraFeedback, since all rewards range from 1 to 5, we set all values to 5 during inference: $p_{1}=p_{2}=p_{3}=p_{4}=5$. For HH-RLHF, there are no unified ranges for each objective as all training values are obtained from reward models. Therefore, we set each preference value to its maximum in the training dataset to enable higher alignment. The values are set to: $p_{1}=4.19$, $p_{2}=3.03$, $p_{3}=1.58$. After calculation, the response with a higher reward value $R_{i}$ is used as the chosen response, and the other response is used as the rejected response for MODPO training. Specifically, the model is trained via the following loss function:
$\mathcal{L}_{\mathrm{CDPO}}={\tiny-\mathbb{E}_{\left(x,c,y_{w},y_{l}\right)\sim\mathcal{D}}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}\mid c,x)}{\pi_{\text{ref }}(y_{w}\mid c,x)}-\beta\log\frac{\pi_{\theta}(y_{l}\mid c,x)}{\pi_{\text{ref }}(y_{l}\mid c,x)}\right)\right]}$ (11)
where $x$ denotes the query, $y_{w},y_{l}$ denote the chosen and rejected prompts, $c$ denotes the corresponding condition, $\pi_{\theta}$ and $\pi_{ref}$ denote the target policy model and the reference model. For implementation, we build our MODPO code based on the OpenRLHF library.
In calculating the GPU hours, we include the training hours for tuning the SFT-based policy model using the CDPO algorithm for HH-RLHF and UltraFeedback datasets. We also include the training hours for fine-tuning the original LLaMA2-Chat-7B model for fair comparisons.
F.3 MORLHF
We use the linear scalarization [27, 20] realization of MORLHF with the KL-divergence regularization. Specifically, the model is trained via the following objective function:
$\mathop{argmax}\limits_{\pi_{\phi}}\mathbb{E}_{q\sim\mathcal{D},y\sim\pi_{\phi}}\left[\omega^{\mathrm{T}}\mathbf{R}(q,y)-\beta log\frac{\pi_{\phi}(y|q)}{\pi_{ref}(y|q)}\right]$ (12)
where $\omega=[\omega_{1},...,\omega_{N}]\ s.t.\sum_{i=1}^{N}\omega_{i}=1,\omega_{i}\geq 0$ is the heuristic target preference vector, $q$, $y$ denote the query and the response, R denotes the reward functions for the target objectives. In implementing the MORLHF algorithm, we first train a reward model for each objective based on random samples from 50% of the training data. Following the reward models we used in building the HH-RLHF dataset, we select the GPT2-large as foundation models for the reward model, and optimize the reward models using the following pair-wise loss functions:
$\mathcal{L}_{rm}=-log(\sigma(R_{c}-R_{r}-margin))$ (13)
where $\sigma$ denotes the Sigmoid function, $R_{c}$ and $R_{r}$ denote the reward output of the chosen response and the rejected response, and $margin$ denotes the margin loss for the corresponding response pairs when multiple responses are ranked, such as in UltraFeedback. Secondly, we fine-tune the LLaMA2-Chat-7B policy model with the highest-ranked responses from the HH-RLHF and UltraFeedback datasets to obtain sub-optimal starting points for RLHF. The SFT training process is formalized as follows:
$\mathop{argmax}\limits_{\pi_{\phi}}\mathbb{E}_{(q,y)\sim\mathcal{D}}\left[P_{\pi_{\phi}}(y|q)\right]$ (14)
where $q$ and $y$ denote the query and its corresponding highest-ranked response. Thirdly, following most works in RLHF, we leverage the PPO algorithm [26] to enable parameterized optimization of the policy model. In linear scalarization, we set $\omega_{1}=\omega_{2}=...=\omega_{N}=\frac{1}{N}$. For implementation, we build our MORLHF code based on the OpenRLHF library.
In calculating the GPU hours, we include the training hours for all reward models in HH-RLHF and UltraFeedback datasets, with a sum of seven reward model training processes. We also include the training hours for fine-tuning the original LLaMA2-Chat-7B model for reaching the sub-optimal starting points. Finally, the PPO training hours for HH-RLHF and UltraFeedback are included in the GPU hours.